OpenStudy (anonymous):

Find the range of the following relation. {(2, 4), (-1, 0), (4, -3), (5, 2)}

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

The "Range" usually refers to the "y" values.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

In a coordinate like (2,4) the 4 is the "y" value.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so you pick the lowest "y" value and the highest "y" value

OpenStudy (anonymous):

{(2, 4), (-1, 0), (4, -3), (5, 2)} In this data set, "-3" is the lowest and "4" is the highest "y" value.

OpenStudy (anonymous):

then you subtract them. 4-(-3) = 4+3 = 7

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes thank you sooo much:D

OpenStudy (anonymous):

In this situation we don't need to worry about all the "x" values (x,y)

OpenStudy (anonymous):

np

OpenStudy (precal):

in this case list all of your y values in stats we do high minus low